OpenStudy (anonymous):

A sum of $2100 is invested in two accounts, one earning 7% simple annual interest and one earning 4%. If the total interest earned for the year is $120, how much money is in each account?

OpenStudy (anonymous):

if you could just help me set the equation up that would be very helpfull

OpenStudy (gw2011):

Let x = 7% investment Let y = 4% investment 0.07x+0.04y=120 x+y=2100 Multiply this equation by 0.04 to find x and you get: 0.07x+0.04Y=120 0.04x+0.04y=84 Subtract this equation from the first one and you get: 0.03x=36 x=1200 Y=2100-1200=900 Therefore, the 7% investment is $1200 and the 4% investment is $900